OpenStudy (anonymous):

Find the exact values of the six trigonometric functions of 165 degrees

Nnesha (nnesha):

more information ?

Nnesha (nnesha):

i guess you have to use sum formula which two angle add up to 165 so \[\huge\rm sin(x+y) = \sin x \cos y + \cos x \sin y\]